OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4

OpenStudy (anonymous):

You can simplify with 3x+3 = -2x + 2 -4

OpenStudy (anonymous):

And it should also be 5 = -5x. Wow it's a slow day XD

OpenStudy (anonymous):

so the answer is -1x

OpenStudy (anonymous):

Well you have to solve for x right? Divide -5 by both sides You will end up with x = -1

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Just saving the world . . . one question at a time.

OpenStudy (anonymous):

whoops i see what u mean i didnt mean write -1x slow day indeed

OpenStudy (anonymous):

3(x+1)=-2(x-1)-4 3x+3=-2x+2-4 3x+2x+3-2+4=0 5x+5=0 5x=-5 x=-5/5 x=-1